A 0.342 T magnetic field in a mass spectrometer causes an isotope of sodium to move in a circular path with a radius of 0.049 m. The charge on an ion is 1.60218 × 10−19 C. If the ions are moving with a speed of 11800 m/s, what is the isotope’s mass? Answer in units of kg.
Answer #1

qvB = mv^2/r

=> 1.602*10^-19*0.342 = m*11800/0.049

=> m = 2.275*10^-25 kg